How many possible samples of size n=8 can be drawn from a population of size 11

1
Expert's answer
2022-02-22T10:46:38-0500

"i)" without replacement

The size of the population is "N=11". If samples of size "n=8" are to be selected from this population without replacement then, the number of possible samples is "\binom{N}{n}=\binom{11}{8}=165".


"ii)" with replacement

The number of possible samples of size "n=8" that can be selected from a population of size "N=11" with replacement is, "N^n=11^8=214358881"